T=gr + gvt for g I divided off r+vt from both sides to place the g alone on right side but then I got stuck
Answer by Aldorozos(172)

Here you are trying to calculate what g is. In other words you are trying to isolate g. For simplicity I rewrite the equation as follow:
gr + gvt = T
Now I factor out g
g(r+vt) = T
Now I divide both sides by (r+vt)
I know that if I divide (r+vt) by itself I get one. Therefore I will have
g(1) = T/(r+vt)
Therefore:
g = T/(r+vt)